In one season, Ana ran 18 races. This was four fewer races than twice the number of races Kelly ran. How many races did Kelly run?
Answer by gwendolyn(128)

Let A be the number of races Ana ran
Let K be the number of races Kelly ran
Ana ran four fewer races than twice the number of races Kelly ran.
So:
A = 2*K - 4 (four fewer means subtract 4)
We already know that Ana ran 18 races, so:
18 = 2K - 4
Add 4 to both sides:
18 + 4 = 2K - 4 + 4
22 = 2K
Divide both sides by 2:
(22)/2 = (2K)/2
11 = K